OpenStudy (bigbosssaint21):

The following table shows the length and width of two rectangles: Length Width Rectangle A 2x + 8 x - 1 Rectangle B 3x + 5 2x - 2 Which expression is the result of the perimeter of rectangle B minus the perimeter of rectangle A? 2x - 4 8x + 10 16x + 20 4x - 8

OpenStudy (bigbosssaint21):

@Whitemonsterbunny17

OpenStudy (bigbosssaint21):

please help!

OpenStudy (anonymous):

perimeter = 2(W + L) Let A = (2x+8)+(x-1) B = (3x+5)+(2x-2) answer = 2B - 2A 2A= (2x+8)+(x-1)+(2x+8)+(x-1) 2A= (2x+8)+(2x+8)+(x-1)+(x-1) 2A= (4x+16)+(2x-2) 2A= (6x+14) 2B= (3x+5)+(2x-2)+(3x+5)+(2x-2) 2B= (6x+10)+(2x-2)+(2x-2) 2B= (10x+6) 2B-2A = (10x+6) - (6x+14) 2B-2A = ???